Could you please help me solve the following equation by using the square root method? 9x^2-30x+25=-1
Thank you very much for any help you can provide.

Answer by stanbon(75887)

solve the following equation by using the square root method? 9x^2-30x+25=-1
----------
(3x-5)^2 = -1
(3x-5)^2+1 = 0
Factor to get:
(3x-5-i)(3x-5+i) = 0
3x-5-i = 0 or 3x-5+i = 0
x=(5+i)/3 or x=(5-i)/3
================
Here's the square root method.
(3x-5)^2 = -1
Take the square root of both sides to get:
3x-5 = i or 3x-5 = -i
3x=5+i or 3x = 5-i
x = (5+i)/3 or x = (5-i)/3
====================
Cheers,
Stan H.